Erol invests R8 350 in an amount that pays simple interest. After six years, the amount that he receives back is R12 859. The simple interest rate on the investment, rounded to two decimal places, is


1
Expert's answer
2021-05-24T17:20:04-0400

"12859 -8350 = 8350 \times r \times 6 \\

4509 = 50100r \\

r = \frac{4509}{50100} = 0.09"